find the integrating factor of (x2y + y2)dx + (y3–x3)dy = 0


1
Expert's answer
2021-02-24T06:02:02-0500

find the integrating factor of (x2y + y2)dx + (y3–x3)dy = 0

Solution:

Let's try find integrating factor from the equation:

"m\cdot(\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x})=N\frac{\partial m}{\partial x}-M\frac{\partial m}{\partial y}" , where

"m(x,y)" - integration factor;

"M(x,y)=x^2y+y^2"

"N(x,y)=y^3-x^3"

"m\cdot(x^2+2y-(-3x^2))=(y^3-x^3)\frac{\partial m}{\partial x}" "-(x^2y+y^2)\frac{\partial m}{\partial y}"

"m\cdot(4x^2+2y)=(y^3-x^3)\frac{\partial m}{\partial x}-(x^2y+y^2)\frac{\partial m}{\partial y}"

This is a partial differential equation and there is no an easy way to solve it except the following three special cases:

1)If the expression:

"\frac{(\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x})}{N}"

is a function of "x" only, then "m" is also a function of "x" only and "m(x)" is given by:

"m(x)=\exp{\left(\int\frac{(\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x})}{N}dx\right)}" .

2) If the expression:

"\frac{(\frac{\partial N}{\partial x}-\frac{\partial M}{\partial y})}{M}"

is a function of "y" only, then "m" is also a function of "y" only and "m(y)" is given by:

"m(y)=\exp{\left(\int\frac{(\frac{\partial N}{\partial x}-\frac{\partial M}{\partial y})}{M}dx\right)}" .

Let's check these cases:

"\frac{(\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x})}{N}=\frac{4x^2+2y}{y^3-x^3}" isn't a function of "x" only .

"\frac{(\frac{\partial N}{\partial x}-\frac{\partial M}{\partial y})}{M}=\frac{-(4x^2+2y)}{x^2y+y^2}" isn't a function of "y" only .

So we can conclude that "m" is a function of "x" and "y" and it cannot be found using the technique described above.

3) If "m(x,y)" can be expressed as "m(x,y)=m(\omega(x,y))" , where "\omega" is known function then integrating factor can be found from the following equation:

"\frac{1}{m}\frac{dm}{d\omega}=\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{N\frac{\partial \omega}{\partial x}-M\frac{\partial \omega}{\partial y}}" (*)

The condition of the task says nothing about the form of the function "m(x,y)". And I couldn't find such a function "\omega(x,y)" to exclude "x" and "y" from (*). So it is highly likely that the problem statement is incomplete or contains an error.